Calculate the total entropy change when each of the following reactions are run at 25°C. Is each reaction spontaneous or nonspontaneous?

CH4(g) + H2O(g) ---> CO(g) + 3H2(g)
Answer #1

Total Entropy change (05°) - AS products - Speedtesta As= (197 + (3X131)) – (186 +189) = 590 - 375 05 = 215 ) it is spontane